bill533:

Translate each equation. Then solve. 1. "The quotient of a number and 4 is -7." 2. "Seven-eights of a number is 63." 3. "The product of a number and -3, increased by 8, is 20." 4. "Eleven subtracted from the quotient of a number and five is -13." 5. "Twenty-seven less than twice a number is -1." 6."The sum of 5% of a number and 9 is 16." 7. "A number subtracted from 16 is -3." 8. "The difference between five-halves of a number and 17 is 48." 9. "Ten more than 32% of a number is 26." 10. "Seven less than a number, divided by 5, is -2." 11. "The quotient of a number increased by 13 and -7 is -4." 12. "The product of -7 and a number squared is -28."

bill533:

@AP @Shadow @dude

AP:

1.) \[n \div 4 = -7\] Solve for n.

bill533:

n = -28

AP:

Yup

bill533:

so what does translate the equation means to you ?

AP:

It means change the word sentence into an equation like I did above ^

bill533:

oh ok continue

AP:

Try translating #2

bill533:

2. 7 divided 63

AP:

It looks like this \[\frac{ 7 }{ 8}n = 63\]

bill533:

n = 72

AP:

You got it.

AP:

3.) \[(n \times -3)+8=20\]

bill533:

n = -4

AP:

Yes 4.)\[(n \div 5) -11=-13\]

bill533:

n = -10

AP:

5.) \[(2 \times n)-27=-1\]

bill533:

n = 13

AP:

6.) 5% can convert to 0.05, so \[(0.05 \times n)+9=16\]

bill533:

n = 140

bill533:

?

AP:

7.) 16 - n = -3

bill533:

n = 19

AP:

8.) \[\frac{ 5 }{ 2 }n - 17 = 58\]

bill533:

n = 30

AP:

9.) 32% = 0.32, so \[(0.32 \times n) +10= 26\]

bill533:

n = 50

AP:

10.) \[\frac{ n-7 }{ 5 }=-2\]

bill533:

-3 never mind

bill533:

right ?

AP:

Yea 11.) \[\frac{ n+13 }{ -7}=-4\]

bill533:

n = 15

AP:

12.) \[-7 \times n^2 = -28\]

bill533:

2 or -2 ?

AP:

Neither

AP:

You can use either Both are correct

bill533:

https://i.imgur.com/D4ybH74.jpg